# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

### v1

@@ 6.3.2 Experiment Results
Edited Instances Base Enhanced Improve Base Enhanced Improve Base Enhanced Improve
GPT-2 (1.5B) 13.6 15.9 16.9% $\uparrow$ 11.6 8.9 23.3% $\downarrow$ 7.0 8.4 20.0% $\uparrow$
GPT-J (6B) 23.1 51.9 124.7% $\uparrow$ 10.1 30.5 202.0% $\uparrow$ 21.3 49.8 133.8% $\uparrow$
LLaMA (7B) 47.7 65.1 36.4% $\uparrow$ 29.6 39.3 32.8% $\uparrow$ 52.4 62.9 20.0% $\uparrow$
Vicuna-v1.5 (7B) 41.3 56.3 36.3% $\uparrow$ 22.7 29.7 30.9% $\uparrow$ 31.6 53.1 68.2% $\uparrow$
LLaMA2 (7B) 36.7 58.7 60.0% $\uparrow$ 17.0 30.3 78.4% $\uparrow$ 29.2 49.1 68.1% $\uparrow$
Falcon (7B) 42.3 61.7 45.7% $\uparrow$ 23.2 31.7 35.7% $\uparrow$ 33.3 48.6 45.7% $\uparrow$
Mistral-v0.1 (7B) 49.0 69.3 41.5% $\uparrow$ 30.0 42.3 41.1% $\uparrow$ 48.7 63.2 29.9% $\uparrow$
Mistral-v0.2 (7B) 44.0 56.3 28.0% $\uparrow$ 23.0 37.7 63.8% $\uparrow$ 32.9 56.2 70.9% $\uparrow$
GPT-J and LLaMA exhibit high fidelity scores across different question types, indicating a strong adherence to the given reasoning paths. Conversely, models such as Vicuna-v1.5, LLaMA2, Falcon, Mistral-v0.1, and Mistral-v0.2 show relatively high mislead accuracy scores with lower fidelity scores. In the experiments, we observe that these models usually rely on their own generated thoughts instead of using incorrect information provided in the CoT. Mistral-v0.2, in particular, demonstrates the lowest fidelity scores and highest misleading accuracy scores, suggesting a potential self-defense ability against false information. The lower fidelity scores of later models may be attributed to their improved training processes on more diverse and high-quality datasets, enabling them to develop a better understanding of context and reasoning. As a result, they are more likely to generate their own correct reasoning paths.
In conclusion, the generated thoughts can be generally viewed as faithful explanations of their output answer. While high fidelity scores generally indicate a model’s adherence to the provided CoT, low fidelity scores do not necessarily imply a lack of faithfulness, especially when the model demonstrates the ability to reject misleading information. Further research on CoT faithfulness and the development of more sophisticated evaluation metrics could contribute to the advancement of interpretable and reliable language models.
Table 6: CoT Faithfulness Evaluation on MQUAKE-CF.
Datasets MQUAKE-CF
Question Type 2-hops 3-hops 4-hops
Edited Instances Correct Mislead Fidelity Correct Mislead Fidelity Correct Mislead Fidelity
@@
6.4 Challenges
@@
Within machine learning, explanation faithfulness refers to the degree to which an explanation accurately reflects the decision-making process of the model it is intended to elucidate (Li et al., 2022c). An explanation is considered as faithful if it causes the model to make the same decision as the original input. In this context, the challenge faced by explainable prompting (e.g., CoT prompt) lies in two aspects: (1) directing language models to generate explanations that are genuinely representative of the models’ internal decision-making processes, and (2) preventing language models from depending on potentially biased CoT templates.
@@
Regarding the first challenge, our case study has revealed that relatively small language models may generate answers that do not align with the provided CoT rationales. Therefore, these rationales do not accurately represent the decision-making process within these models. Some efforts have been made to bolster the CoT capabilities of smaller language models by implementing instruction tuning with CoT rationales (Kim et al., 2023b; Ho et al., 2022). These methods can help improve the explanation faithfulness of CoT for small language models, thereby addressing this issue to some extent. Nevertheless, it remains a challenging problem of how to ensure the generated explanations (i.e., “what the model says”) are faithful to the internal mechanism (i.e., “what the model thinks”) of language models.
Regarding the second challenge, recent research shows that explanations in the CoT can be heavily influenced by the introduction of biasing prompt templates into model input (Turpin et al., 2024). This is because existing CoT requires carefully designed templates to prompt language models to produce explanations. If incorrect or biased information is encoded in such templates, the generated explanations could be misleading. Recently, Wang & Zhou (2024a) propose a novel decoding strategy to implement CoT with prompting, which could mitigate this issue. However, how to effectively help language models get rid of the template reliance still remains to be underexplored.
7 LLM Enhancement via Knowledge-Augmented Prompting
@@
Enhancing models with external knowledge can significantly improve the control and interpretability of decision-making processes. While LLMs acquire extensive knowledge through pre-training on web-scale data, this knowledge is embedded implicitly within the model parameters, making it challenging to explain or control how this knowledge is utilized during inference. Additionally, LLMs may not always encompass the unique knowledge specific to certain domains, nor keep pace with the constantly evolving information in the world. To address these limitations, this section discusses Retrieval-Augmented Generation (RAG) for the explicit integration of external knowledge into the mechanism of LLM-based AI systems, aiming to yield more interpretable predictions.
7.1 Preliminaries: Retrieval-Augmented Generation
By fetching relevant information from external databases or the internet, RAG ensures that LLM outputs are accurate and up-to-date. It addresses LLMs’ limitation of relying on fixed and potentially outdated knowledge bases. RAG operates in two steps: (1) Retrieval: It locates and fetches pertinent information from an external source based on the user’s query; (2) Generation: It incorporates this information into the model’s generated response. Given an input query $x$ and the desired output $y$, the objective function of RAG can be formulated as (Guu et al., 2020):
$\max_{\phi,\theta}\,\log p(y|x)=\max_{\phi,\theta}\,\log\sum_{z\in\mathcal{K}}p_{\phi}(y|x,z)\cdot p_{\theta}(z|x),$ (7)
@@ 7.1 Preliminaries: Retrieval-Augmented Generation
where $z$ stands for the external knowledge retrieved from a knowledge base $\mathcal{K}$. Thus, the target distribution is jointly modeled by a knowledge retriever $p_{\theta}(z|x)$ and an answer reasoning module $p_{\phi}(y|x,z)$. The knowledge $z$ serves as a latent variable. An RAG model is trained to optimize the parameters, so that it learns to retrieve relevant knowledge $z$ and to produce correct answers $y$ based on $z$ and $x$. As LLMs possess stronger text comprehension and reasoning abilities, they can directly serve as the reasoning module $p_{\phi}$ without further training. In this case, RAG can be treated as a data-centric problem:
$\max_{z\in\mathcal{K}}\,\log p(y|z,x)=\max_{z\in\mathcal{K}}\,\frac{p(z|x,y)}{p(z|x)}p(y|x),$ (8)
@@
where the goal is to find appropriate knowledge that supports the desired output. The interpretability of RAG-based models comes from the information in $z$: (1) $z$ usually elucidates or supplements the task-specific information in $x$; (2) $z$ could explain the generation of output $y$. Unlike other deep models that directly estimate $p(y|x)$ in an end-to-end manner, where the decision process is not comprehensible, the RAG process provides justification or rationale $z$ that supports the result.
Existing Retrieval-Augmented Generation (RAG) approaches can be categorized based on when they integrate external knowledge into the model’s workflow. The first category incorporates external knowledge at the inference stage. For instance, Karpukhin et al. (2020) employ dense vectors to identify related documents or text passages, enhancing the data retrieval step of RAG. Similarly, Lewis et al. (2020) refine the data retrieval process to ensure only the most pertinent information influences the model’s output. The second category integrates external knowledge during the model tuning stage. Some representative approaches include Guu et al. (2020); Borgeaud et al. (2022); Nakano et al. (2021). Generally, these methods embed a retrieval mechanism into the model’s training phase, enabling the model to utilize external data more efficiently from the outset.
7.2 Enhancing Decision-Making Control with Explicit Knowledge
The incorporation of explicit external knowledge through RAG enhances the precision and controllability of decision-making in LLMs. This method leverages real-time information from external databases to produce responses that are not only accurate but also tailored to the specific requirements of each query. Below, we explore the mechanisms by which RAG achieves a more controllable and directed content generation process, with references to key papers that have contributed to these advancements.
7.2.1 Reducing Hallucinations in Response
@@
“Hallucination" in the context of LLMs refers to instances where these models generate information that, while coherent and contextually appropriate, is not based on factual accuracy or real-world evidence (Huang et al., 2023c). This issue can lead to the production of misleading or entirely fabricated content, posing a significant challenge to the reliability and trustworthiness of LLMs’ outputs. RAG offers a powerful solution to mitigate the problem of hallucinations in LLMs. By actively incorporating up-to-date, verified external knowledge at the point of generating responses, RAG ensures that the information produced by the model is anchored in reality. This process significantly enhances the factual basis of the model’s outputs, thereby reducing the occurrence of hallucinations. Shuster et al. (2021) applies neural-retrieval-in-the-loop architectures to knowledge-grounded dialogue, which significantly reduces factual inaccuracies in chatbots, as confirmed by human evaluations. Siriwardhana et al. (2023) introduces RAG-end2end, which joint trains retriever and generator components together. Their method demonstrates notable performance improvements across specialized domains like healthcare and news while reducing knowledge hallucination.
7.2.2 Dynamic Responses to Knowledge Updating
@@
RAG empowers LLMs with the ability to incorporate the most current information, keeping their decision-making processes aligned with the latest developments. This feature is especially vital in fast-evolving fields such as medicine and technology, where the need for timely and accurate information is paramount (Meng et al., 2022b). For example, research by (Izacard & Grave, 2020) demonstrates significant enhancements in output relevance and accuracy through real-time information retrieval. Similarly, Han et al. (2023) suggest using retrieved factual data to correct and update the knowledge within pre-trained LLMs efficiently. Additionally, Wang et al. (2023e) introduce a method for integrating newly retrieved knowledge from a multilingual database directly into the model prompts, facilitating updates in a multilingual context.
7.2.3 Domain-specific Customization
@@
RAG enhances LLMs by incorporating knowledge from specialized sources, enabling the creation of models tailored to specific domains. Research by Guu et al. (2020) illustrates how integrating databases specific to certain fields into the retrieval process can empower models to deliver expert-level responses, boosting their effectiveness in both professional and academic contexts. Shi et al. (2023) have applied this concept in the medical domain with MedEdit, utilizing an in-context learning strategy to merge relevant medical knowledge into query prompts for more accurate medical advice. Moreover, recent research finds that LLMs struggle to capture specific knowledge that is not widely discussed in the pre-training data. Specifically, Mallen et al. (2023) observe that LLMs often fail to learn long-tail factual knowledge with relatively low popularity, finding that simply increasing model size does not significantly enhance the recall of such information. However, they note that retrieval-augmented LLMs surpass much larger models in accuracy, particularly for questions on well-known subjects, suggesting that this method can effectively bridge knowledge gaps. Similarly, Kandpal et al. (2023) highlights LLMs’ challenges with acquiring rare knowledge and proposes that retrieval augmentation offers a viable solution, minimizing reliance on extensive pre-training for capturing nuanced, less common information.
7.3 Challenges
@@
We discuss the challenges in RAG that are relevant to its explainability aspects: (1) In the retrieval stage $p_{\theta}(z|x)$, does the retrieved information $z$ always elucidate the task-specific information contained in the input $x$? (2) In the generation stage $p_{\phi}(y|x,z)$, does $z$ effectively serve as an explanation for the generation of output $y$? Please note that our goal is not to exhaustively discuss all the limitations of RAG in this paper as RAG itself is a broad topic in NLP research. For a more detailed examination of the broader limitations of RAG, we direct readers to other reviews (Gao et al., 2023).
7.3.1 Retrieval Accuracy Bottlenecks
@@
Existent RAG methods typically rely on similarity search to pinpoint relevant information (Lewis et al., 2020; Gao et al., 2023), which represents a substantial improvement over basic keyword searches (Robertson et al., 2009). However, these methods may struggle with complex queries that demand deeper comprehension and nuanced reasoning. The recent “lost-in-the-middle” phenomenon (Liu et al., 2024) has revealed that an ineffective retrieval can result in the accumulation of extraneous or conflicting information, negatively affecting the generation quality. To address this challenge, recent RAG approaches have integrated adaptive learning processes (Asai et al., 2023). This advancement enables the retrieval system to refine their performance over time through feedback, adapting to evolving language use and information updates, ensuring their responses remain relevant and accurate. Nonetheless, efficiently handling intricate and multi-hop questions remains a significant challenge, highlighting the need for ongoing research to enhance the capabilities of RAG systems.
7.3.2 Controllable Generation Bottlenecks
@@
In-context learning stands out as the premier method for incorporating external knowledge to boost the capabilities of LLMs such as GPT-4 (Asai et al., 2023; Gao et al., 2023). Despite its effectiveness, there’s no surefire way to ensure that these models consistently leverage the provided external knowledge within the prompts for their decision-making processes. In practice, to achieve thorough coverage, commonly used dense retrieval usually returns a large volume of content, including both relevant and redundant information to the input question. Unfortunately, redundant information in the model prompt raises the computational cost and can mislead LLMs to generate incorrect answers. Recent research shows the retrieved information can degrade the question-answering task performance (Yoran et al., 2023; Petroni et al., 2020; Li et al., 2022a). Some recent work proposes to fine-tune the LLM to improve resilience to noise and reduce hallucinations. However, such approach still cannot prevent oversized retrieval information decrease the system interpretability (Yoran et al., 2023; Xu et al., 2023b). The challenge of optimizing the use of external explanations to achieve more precise and controlled decision-making in LLMs is an ongoing issue that has yet to be fully addressed.
8 Training Data Augmentation with Explanation
@@
This section explores the generation of synthetic data from explanations using large language models, a technique poised to enhance various machine learning tasks. In machine learning, limited data availability often constrains model performance, presenting a significant challenge across many domains. A viable solution is data augmentation, where LLMs, with their robust generative capabilities, can be utilized for text data augmentation (Whitehouse et al., 2023), such as transforming existing text samples into new variants (Dai et al., 2023). Nevertheless, there are several challenges to be tackled for effective text augmentation. First, for utility, the generated samples need to exhibit diversity compared to the original data. Second, these samples should be exhibit useful patterns relevant to the downstream tasks. To address these challenges, explanation methods offer a valuable tool to guide data augmentation by providing supplemental contexts and useful rationales (Carton et al., 2021). Using LLMs for explanation-guided data augmentation is a nascent but promising field. In this section, we aim to outline feasible frameworks and discuss potential applications, offering directions for future research in this field.
@@
Explanations can be particularly beneficial in data augmentation within two scenarios. In the first scenario, explanations are used to delineate desired model behaviors or to identify existing deficiencies, which effectively guides the data augmentation process of LLMs. The second scenario involves employing LLMs to directly produce explanatory texts, which serve as supplementary information to enrich the dataset.
8.1 Explanation-guided Data Augmentation for Mitigating Shortcuts

### v2

@@ 6.2.2. Experiment Results
Edited Instances C M F C M F C M F
@@
C: CoT; M: Misleaded CoT; F: Fidelity Score.
GPT-J and LLaMA exhibit high fidelity scores across different question types, indicating a strong adherence to the given reasoning paths. Conversely, other models show relatively high mislead accuracy scores with lower fidelity scores. In the experiments, we observe that these models usually rely on their own generated thoughts instead of using incorrect information provided in the CoT. Mistral-v0.2, in particular, demonstrates the lowest fidelity scores and highest misleading accuracy scores, suggesting a potential self-defense ability against false information. The lower fidelity scores of later models may be attributed to their improved training processes on more diverse and high-quality datasets, enabling them to develop a better understanding of context and reasoning. As a result, they are more likely to generate their own correct reasoning paths.
6.3. Challenges
@@
An explanation is considered as faithful if it causes the model to make the same decision as the original input (Li et al., 2022c). In this context, the challenge faced by explainable prompting (e.g., CoT prompt) lies in two aspects: (1) directing language models to generate explanations that are genuinely representative of the models’ internal decision-making processes, and (2) preventing language models from depending on potentially biased CoT templates.
@@
Regarding the first challenge, our case study has revealed that relatively small language models may generate answers that do not align with the provided CoT rationales. Therefore, these rationales do not accurately represent the decision-making process within these models. Some efforts have been made to bolster the CoT capabilities of smaller language models by implementing instruction tuning with CoT rationales (Kim et al., 2023a; Ho et al., 2022). Nevertheless, it remains a challenging problem of how to ensure the generated explanations (i.e., “what the model says”) are faithful to the internal mechanism (i.e., “what the model thinks”) of language models. Regarding the second challenge, recent research shows that explanations in the CoT can be heavily influenced by the introduction of biasing prompt templates into model input (Turpin et al., 2024). If incorrect or biased information is encoded in such templates, the generated explanations could be misleading. Recently, (Wang and Zhou, 2024) propose a novel decoding strategy to implement CoT with prompting, which could mitigate this issue. However, how to effectively help language models get rid of the template reliance still remains underexplored.
7. Effective LLM Inference via Knowledge-Augmented Prompting
@@
Enhancing models with external knowledge can significantly improve the control and interpretability of decision-making processes. While LLMs acquire extensive knowledge through pre-training on web-scale data, this knowledge is embedded implicitly within the model parameters, making it challenging to explain or control how this knowledge is utilized during inference. To address these limitations, this section discusses Retrieval-Augmented Generation (RAG) for the explicit integration of external knowledge into LLMs, aiming to yield more interpretable predictions.
7.1. Retrieval-Augmented Generation with Knowledge
@@
Generally, RAG operates in two steps: (1) Retrieval: It locates and fetches pertinent information from an external source based on the user’s query; (2) Generation: It incorporates this information into the model’s generated response. Given an input query $x$ and the desired output $y$, the objective function of RAG can be formulated as (Guu et al., 2020): $\max_{\phi,\theta}\,\log p(y|x)=\max_{\phi,\theta}\,\log\sum_{z\in\mathcal{K}}p_{\phi}(y|x,z)\cdot p_{\theta}(z|x),$ where $z$ stands for the external knowledge retrieved from a knowledge base $\mathcal{K}$. Thus, the target distribution is jointly modeled by a knowledge retriever $p_{\theta}(z|x)$ and an answer reasoning module $p_{\phi}(y|x,z)$. The knowledge $z$ serves as a latent variable. An RAG model is trained to optimize the parameters, so that it learns to retrieve relevant knowledge $z$ and to produce correct answers $y$ based on $z$ and $x$. As LLMs possess stronger text comprehension and reasoning abilities, they can directly serve as the reasoning module $p_{\phi}$ without further training. In this case, RAG can be treated as a data-centric problem:
(12) $\max_{z\in\mathcal{K}}\,\log p(y|z,x)=\max_{z\in\mathcal{K}}\,\frac{p(z|x,y)}{p(z|x)}p(y|x),$
@@
where the goal is to find appropriate knowledge that supports the desired output. The interpretability of RAG-based models comes from the information in $z$: (1) $z$ usually elucidates or supplements the task-specific information in $x$; (2) $z$ could explain the generation of output $y$. Unlike other models that estimate $p(y|x)$ in an end-to-end manner, where the decision process is not comprehensible, the RAG process provides justification or rationale $z$ that supports the result.
7.2. Enhancing Decision-Making Control with Explicit Knowledge
7.2.1. Reducing Hallucinations in Response
@@
“Hallucination” in the context of LLMs refers to instances where these models generate information that, while coherent and contextually appropriate, is not based on factual accuracy or real-world evidence (Huang et al., 2023d). This issue can lead to the production of misleading or entirely fabricated content, posing a significant challenge to the reliability and trustworthiness of LLMs’ outputs. RAG offers a powerful solution to mitigate the problem of hallucinations in LLMs. By actively incorporating up-to-date, verified external knowledge at the point of generating responses, RAG ensures that the information produced by the model is anchored in reality. This process significantly enhances the factual basis of the model’s outputs, thereby reducing the occurrence of hallucinations. (Shuster et al., 2021) applies neural-retrieval-in-the-loop architectures to knowledge-grounded dialogue, which significantly reduces factual inaccuracies in chatbots, as confirmed by human evaluations. (Siriwardhana et al., 2023) introduces RAG-end2end, which joint trains retriever and generator components together. Their method demonstrates notable performance improvements across specialized domains like healthcare and news while reducing knowledge hallucination.
7.2.2. Dynamic Responses to Knowledge Updating
@@
RAG empowers LLMs by incorporating the most up-to-date information, keeping their decision-making aligned with the latest developments. This feature is especially vital in fast-evolving fields such as medicine and technology, where the need for timely and accurate information is paramount (Meng et al., 2022b). For example, research by (Izacard and Grave, 2020) demonstrates significant enhancements in output relevance and accuracy through real-time information retrieval. Similarly, (Han et al., 2023) suggest using retrieved factual data to correct and update the knowledge within pre-trained LLMs efficiently. Additionally, (Wang et al., 2023b) introduce a method for integrating newly retrieved knowledge from a multilingual database directly into the model prompts, facilitating updates in a multilingual context.
7.2.3. Domain-specific Customization
@@
RAG enhances LLMs by incorporating knowledge from specialized sources, enabling the creation of models tailored to specific domains. Research by (Guu et al., 2020) illustrates how integrating databases specific to certain fields into the retrieval process can empower models to deliver expert-level responses, boosting their effectiveness in both professional and academic contexts. (Shi et al., 2023) have applied this concept in the medical domain with MedEdit, utilizing an in-context learning strategy to merge relevant medical knowledge into query prompts for more accurate medical advice. Moreover, recent research finds that LLMs struggle to capture specific knowledge that is not widely discussed in the pre-training data. Specifically, (Mallen et al., 2023) observe that LLMs often fail to learn long-tail factual knowledge with relatively low popularity, finding that simply increasing model size does not significantly enhance the recall of such information. However, they note that retrieval-augmented LLMs surpass much larger models in accuracy. Similarly, (Kandpal et al., 2023) highlights LLMs’ challenges with acquiring rare knowledge and proposes that retrieval augmentation offers a viable solution, minimizing reliance on extensive pre-training for capturing nuanced, less common information.
7.3. Challenges
@@
We discuss the challenges in RAG that are relevant to its explainability aspects: (1) In the retrieval stage $p_{\theta}(z|x)$, does the retrieved information $z$ always elucidate the task-specific information contained in the input $x$? (2) In the generation stage $p_{\phi}(y|x,z)$, does $z$ effectively serve as an explanation for the generation of output $y$?
7.3.1. Retrieval Accuracy Bottlenecks
@@
Existent RAG methods typically rely on similarity search to pinpoint relevant information (Lewis et al., 2020; Gao et al., 2023b), which represents a substantial improvement over basic keyword searches (Robertson et al., 2009). However, these methods may struggle with complex queries that demand deeper comprehension and nuanced reasoning. The recent “lost-in-the-middle” phenomenon (Liu et al., 2024a) has revealed that an ineffective retrieval can result in the accumulation of extraneous or conflicting information, negatively affecting the generation quality. Efficiently handling intricate and multi-hop questions remains a significant challenge, highlighting the need for ongoing research to enhance the capabilities of RAG systems.
7.3.2. Controllable Generation Bottlenecks
@@
In-context learning stands out as the premier method for incorporating external knowledge to boost the capabilities of LLMs such as GPT-4 (Asai et al., 2023b; Gao et al., 2023b). Despite its effectiveness, there’s no surefire way to ensure that these models consistently leverage the provided external knowledge within the prompts for their decision-making processes (Wu et al., 2024b; Yoran et al., 2023; Petroni et al., 2020; Li et al., 2022b). The challenge of optimizing the use of external explanations to achieve more precise and controlled decision-making in LLMs is an ongoing issue that has yet to be fully addressed.
8. Training Data Augmentation with Explanation
@@
This section explores the generation of synthetic data from explanations using LLMs, a technique poised to enhance various machine learning tasks. In machine learning, limited data availability often leads to data imbalance and scarcity, which in turn constrains model performance in many domains. A viable solution is data augmentation, where LLMs, with advanced generative capabilities, can be utilized (Whitehouse et al., 2023; Tan et al., 2024). Nevertheless, several challenges exist for effective text augmentation. First, for utility, the generated samples should exhibit diversity compared to the original data. Second, these samples should exhibit useful patterns relevant to downstream tasks. To this end, explanation tools could guide augmentations by providing supplemental context and rationales (Carton et al., 2021).
@@
Explanations can be particularly beneficial in two scenarios. In the first scenario, explanation tools are used to to identify existing deficiencies, which effectively guides the data augmentation. The augmented data include samples and related label annotations. The second scenario employs LLMs to directly produce explanatory texts, such as rationales, as supplementary information to enrich the dataset.
8.1. Label-based Data Augmentation
For learning tasks, the training dataset includes a large volume of data points, defined as $\mathcal{D}=\{(x_{1},y_{1}),\cdots,(x_{n},y_{n})\}$. We categorize learning tasks into classification tasks and generation tasks.